A penny moving at speed v slides off the horizontal surface of a coffee table a vertical distance y from the floor.
(a) Show that the penny lands a distance
1
Expert's answer
2020-05-04T12:47:53-0400

Equations of motion of a penny are as follows


"x(t)=v_0t\\
y(t)=H-\frac{gt^2}{2}"

When penny lands "y=0", hence


"t^*=\sqrt{\frac{2H}{g}}"

"L=x(t^*)=v_0\sqrt{\frac{2H}{g}}"